Miss Sunni manages a herbicide plant. She wants a solution of 45% herbicide to be mixed with a solution of 24% herbicide to form 42 liters of a 34% solution. How many liters of the 45% solution must Miss Sunni use?
I tried using the chart method with this and not sure what to do with the 42 liters?
Thanks, any help would be appreciated!!

Answer by Paul(988)

45x+24(42-x)=34(42)
45x+1008-24x=1428
21x=420
x=20
Hence, for 45% solution there should be a miture of 20L solution